OpenStudy (anonymous):

1.Blair is five times as old as Alex. If the difference in ages is 48 years, how old is Alex? A.5 B.10 C.12 D.15

OpenStudy (anonymous):

thats a hard one

OpenStudy (butterflydreamer):

Let B = Blair, A = Alex. So the first line of your question tells you that "Blair is 5 times as old as Alex" This means that B = 5A. Next the question tells you "The difference in ages if 48 years old". So this means Blair - Alex = 48. Therefore: B - A = 48 We know that B = 5A so then the "B - A = 48" equation becomes: 5A - A = 48 Now what does 5A - A equal to?

OpenStudy (anonymous):

Oh I was trying to build by the simultaneous equation too I got till the part of 5A - A = 48 didn't think of B = 5A. good job @butterflydreamer :)

OpenStudy (anonymous):

*to build the

OpenStudy (butterflydreamer):

hahaha thanks :) @nopen